Grade 11IIT JEE Entrance Exam

show that the equations (a) x=2+t ,y=3+2t and (b)x=3-5t,y5-10t represent the same line.

sneha
10 Years agoGrade 11

1 Answer

Shantanu Singh
10 Years ago
    (a)      x=2+t 
               t=x-2------(1)
              y=3+2t 
               Substituting the value of ‘t’ from equation (1)
              y=2x-1-------(A)
 
      (b).     x=3-5t
                (3-x)/5=t------(2)
                y=5-10t
                Substituting the value of ‘t’ from equation (2)
                y=2x-1--------(B)
 
                (A)=(B), hence proved